Find the minimal polynomial of - 3+ 53 over . Make sure to completely justify all your claims.
Answer #1

et. N= -3+33 (2+3)=33 cube both side (n+3)3 W3+aztan(n+3)= 3 x² eqre²+278 +24=0 34243 This polynomial is irreduriale oka is b